### v1

@@ Abstract
We explore a capability of evolution strategies to train an agent with its policy based on a transformer architecture in a reinforcement learning setting. We performed experiments using OpenAI’s highly parallelizable evolution strategy to train Decision Transformer in Humanoid locomotion environment and in the environment of Atari games, testing the ability of this black-box optimization technique to train even such relatively large and complicated models (compared to those previously tested in the literature). We also proposed a method to aid the training by first pretraining the model before using the OpenAI-ES to train it further, and tested its effectiveness. The examined evolution strategy proved to be, in general, capable of achieving strong results and managed to obtain high-performing agents. Therefore, the pretraining was shown to be unnecessary; yet still, it helped us observe and formulate several further insights.
@@ I Introduction
The problem of reinforcement learning is considered one of the most difficult in the field of machine learning. There are many approaches to solving it [1]. Some are based on computing a gradient to optimize the objective, some are derivative-free. One such class of general derivative-free optimization algorithms are evolutionary algorithms [2]. Their subclass of evolution strategies [3] has been proved to be a viable alternative to gradient approaches for the (deep) reinforcement learning [4]. Although the gradient approaches generally have better sample utilization, the evolution strategies are greatly parallelizable. Moreover, evolution strategies have better exploration of possible solutions and the agents trained using these methods are usually more diverse than those trained by the gradient-based algorithms. They can even incorporate techniques that vastly improve the exploration even more, such as searching for novelty instead of (novelty search [5, 6]), or in addition to (quality-diversity [7]) just seeking a better performance. A fairly simple, yet highly effective example of such algorithms for reinforcement learning is OpenAI-ES [8].
@@
We decided to subject the combination of the evolution strategies and the Decision Transformers to experiments, and test the ability of derivative-free algorithms to train this more complicated and bigger – compared to the simple feed-forward models that had been experimented with in the literature so far – transformer architecture. It might be the case, that for large and complicated models it may be hard, or even almost impossible, to be trained from scratch using an evolutionary approach. Therefore, we wanted to experiment with first pretraining the agent using a supervised learning of sequence prediction on data generated by a smaller, potentially weaker model, which can be trained using any arbitrary reinforcement learning method.
In the following section, we present the background for our experiments. In Section III, we introduce our experiments and show their results, whereas in Section IV, we discuss those results. We then conclude the paper in the last section.
@@ II-A Evolution Strategies
Evolutionary algorithms are a large and quite a successful family of black-box derivative-free optimization algorithms. They consist of various metaheuristics inspired by nature. One subclass of such nature-inspired optimization methods are evolution strategies. Introduced as a tool for dealing with high-dimensional continuous-valued domains [3], evolution strategies work with a population of individuals (real-valued vectors). In each generation (iteration / population in the iteration), they derive a new set of individuals by somehow mutating (pertrubating) the original population; the new set is then evaluated (with respect to a given objective function), and a new generation is formed based on these new evaluated individuals taking into account their fitness (objective function value).
In order to use an evolution strategy as a reinforcement learning algorithm, we use agents represented by a neural network, or, more specifically, by a real-valued vector of the network weights, as the individuals for the algorithm. Then the only thing needed is to set the fitness of each individual as the mean return of the individual agent from several episodes. Various evolution strategy algorithms were proposed for this purpose [4].
@@
However, evolution strategies also bring a lot of advantages with them. Many of the evolution strategy algorithms are easily parallelizable, and much of the research in the area has been focused on this aspect, yielding us algorithms with a linear improvement of performance when more computing power is used [8]. And because evolution strategies are derivative-free algorithms, we can optimize not just classical smooth neural networks, but our models can even contain some discrete subfunctions or be otherwise non-differentiable.
In this paper, we will be working with OpenAI-ES [8], which is a representative of Natural evolution strategies [12]. The population is represented by a distribution over the agent’s (neural network’s) parameters, the distribution being a Gaussian, whose mean value is the candidate solution to the given problem and from which the offsprings are sampled each generation. These are then evaluated, and their evaluation is used to update the parameters (in our case only the mean value) of the distribution so that we obtain a higher expected fitness for the future samples. This update is performed using an approximation of natural gradient (whence the name Natural evolution strategies). In our case, this is obtained by renormalizing (rescaling) the update with respect to uncertainty. (In general, we would have to compute an inverse of a so-called Fisher information matrix, which would then be used to multiply the gradient estimate by. However, as demonstrated, e.g., in a paper by Pierrot, Perrin-Gilbert and Sigaud [13], in our case, when we derive the parameter updates from the Gaussian distribution with the same variance for each parameter, dividing by the variance (the rescaling w.r.t. uncertainty) achieves approximately the same result.) The algorithm is also designed in such a way that it is highly parallelizable with interprocess communication kept at bare minimum. For more details or for a discussion of the design choices, we refer our readers to the original paper [8].
@@ II-B Transformers
The transformers interlace classical feed-forward layers with self-attention layers. (In practice, there are also residual connections and layer normalizations to stabilize the training.) And it is the self-attention to which the transformers owe their success. For each element of an input sequence, the self-attention constructs a "key", a "query", and a "value" using fully connected layers of neurons. Then, to compute an $i$-th element of an output sequence on the same position, it takes a combination of all the values, each weighted proportionally to the product of the query belonging to the given ($i$-th) position and a key corresponding to the value – hence combining the information from the whole input sequence to produce every single element of the output sequence, as shown by the following equation.
@@
Its main idea is that we want the agent’s policy to produce an action based on not just the last observation, but the whole history (or the part which squeezes into the context window) of past observations and undertaken actions. And to have some way to affect the agent’s performance, we add a conditioning on a return-to-go, which is a return we want to obtain from a given step until the end of the episode.
@@ III Experiments
From what we saw in the previous section, the evolution strategies belong to the state-of-the-art algorithms for solving the reinforcement learning problem, whereas the transformers embody a new possibility for how to approach this problem, as well as a potentially strong architecture of the neural network representing the agent’s policy. Therefore, it is just a logical next step to try and combine these two concepts, to test the capability of the transformers to be trained by evolution strategies in the setting of reinforcement learning.
@@
We provided our own implementation of the OpenAI-ES  [^1] Our code, together with the data gathered during all the experiments can be found at the following address: https://github.com/Mafi412/Evolution-Strategies-and-Decision-Transformers and began by a replication experiment of the original paper, which also serves as a correctness check for our code. Here, we tested the algorithm on a classical feed-forward network in the MuJoCo [15] Humanoid environment using OpenAI Gym [16], whereupon we followed up by proceeding to test the performance of the evolution strategy with the Decision Transformer in the same environment. We then examined an idea of pretraining the transformer by a behavior cloning of some smaller and easily trained – yet possibly weaker – model before training the transformer by the evolution strategy. The last experiments we present are studying the particulars of training a larger model in a different environment with a higher-dimensional image input, which is Hero, one of the Atari games [17].
@@
In all the experiments with Decision Transformers, we used the same hyperparameter values for the model as used in the original paper [10] for the respective environments. For comparison, the model sizes are 166 144 parameters for the feed-forward model from the original OpenAI-ES paper [8], 825 098 parameters for the Decision Transformer for MuJoCo Humanoid environment and 2 486 272 parameters for Atari games environment.
For all experiments, we conducted ten runs of the training and aggregated the results in our plots. For every run of each of the experiments, 300 workers were used utilizing 301 CPU cores (with one being a master handling synchronization, evaluation, and saving the agent). The desired returns passed to all the Decision Transformer models at the beginning of each episode were 7000 for the Humanoid Decision Transformer and 8000 for the Atari Decision Transformer.
@@
(a) Evaluation results
(b) Population fitnesses
Figure 2: OpenAI-ES used on a simple feed-forward model for a MuJoCo Humanoid simulation in ten runs of the experiment. In Figure 2(a), we can see the results of evaluations of the current solution model after every iteration of the algorithm. Figure 2(b) shows the fitness values of the whole populations throughout the iterations (mean and percentile interval with standard width, ranging from 2.5 to 97.5 percentiles).
Our Implementation
Mainly due to different computational resources available to us – we needed to use the MPI framework instead of EC2 (AWS) – as well as the need to interchange the models being trained and the environments for the evaluation more easily and with a greater degree of freedom, we decided to create our own implementation of OpenAI-ES. Yet, in the process of creating this implementation of ours we noticed some inconsistencies between the original paper introducing the algorithm [8] and their provided code. We decided to adhere more to the paper when these dissimilarities occured and here we state our reasoning for the two most significant differences.
First and foremost, in the paper [8], the use of a weight decay is mentioned as a form of keeping the effect of newly added updates still significant enough. After every update, we decay the weights of our model, which is the distribution mean. Therefore, it would not happen that there could be weights grown to such sizes that a newly added noise practically does not change the function of the model. This reasoning tells us why we want to use preciselly the weight decay. Nevertheless, in the original implementation L2-regularization is used instead. This may be all right for some optimizers, as for, e.g., SGD or SGD with momentum it is virtually the same as the weight decay (just rescaled by the learning rate). However, as mentioned in paper Decoupled Weight Decay Regularization [18], it is not the same, for example, for ADAM optimizer, which is nonetheless exactly the one used in the original experiments. Our implementation, on the other hand, remains true to the original paper and uses proper weight decay. Nevertheless, we employ SGD with momentum in our experiments, as it demonstrated superior efficacy in our setting during our experiments.
Secondly, during the update of the distribution parameters, a gradient estimate should be normalized with respect to the uncertainty using a division by the standard deviation, so that it becomes a natural gradient estimate. Despite that, in the original OpenAI code, this is not done. There, this is probably hidden in the value of a learning rate (which, as a result, differs from ours). However, then the learning rate and the noise deviation are unnecessarily coupled hyperparameters.
In addition to the implementation of OpenAI-ES, we made a few slight modifications even in the Decision Transformer code. The most significant one concerns keeping data from the past to be used in future inputs. In the original code, the authors do not clip sequences already longer than the context window; those are hence allowed to grow unconstrained in the memory. The new data are always just appended to the list, keeping the whole history, and the whole list is passed to an inference function, which only then clips the sequence to the required length (without modifying the data stored) and passes it to the transformer. This proved to be no problem in MuJoCo environment, but for some reason leads to a severe slowdown in the Atari environment. Thus, we keep the stored sequence cropped to match the context length in the Atari version of Decision Transformer.
(a) Evaluation results
(b) Evaluation results with half the population
Figure 3: OpenAI-ES used on a Decision Transformer model for a MuJoCo Humanoid simulation in ten runs of the experiment. Figure 3(a) shows us the evaluation results of the current solution in each generation of the evolution strategy, whereas in Figure 3(b) we can see the evaluations for the experiment using just half the population.
(a) Evaluation results
(b) Evaluation results with half the population
Figure 4: OpenAI-ES used on a pretrained Decision Transformer model for a MuJoCo Humanoid simulation in ten runs of the experiment. In Figure 4(a), the evaluation results of the current solution of each iteration are shown, and Figure 4(b) presents the evaluation results for this experiment using only half the population.
III-A Feed-Forward - Humanoid

### v2

@@ Abstract
We explore the capability of evolution strategies to train an agent with a policy based on a transformer architecture in a reinforcement learning setting. We performed experiments using OpenAI’s highly parallelizable evolution strategy to train Decision Transformer in the MuJoCo Humanoid locomotion environment and in the environment of Atari games, testing the ability of this black-box optimization technique to train even such relatively large and complicated models (compared to those previously tested in the literature). The examined evolution strategy proved to be, in general, capable of achieving strong results and managed to produce high-performing agents, showcasing evolution’s ability to tackle the training of even such complex models.
@@ I Introduction
The problem of reinforcement learning is considered one of the most difficult in the field of machine learning. There are many approaches to solving it [1]. Some are based on computing a gradient to optimize the objective, some are derivative-free. One such class of general derivative-free optimization algorithms are evolutionary algorithms [2]. Their subclass of evolution strategies [3] has been proved to be a viable alternative to gradient approaches for the (deep) reinforcement learning [4]. Although the gradient approaches generally have better sample utilization, the evolution strategies are greatly parallelizable. Moreover, evolution strategies have better exploration of possible solutions and the agents trained using these methods are usually more diverse than those trained by the gradient-based algorithms. They can even incorporate techniques that vastly improve the exploration even more, such as searching for novelty instead of, or in addition to just seeking better performance. This yields novelty search [5, 6] or quality-diversity [7] algorithms. An example of such algorithm for reinforcement learning, that is fairly simple, yet highly efficient, is OpenAI-ES [8].
@@
We decided to subject the combination of the evolution strategies and the Decision Transformers to experiments, and test the ability of derivative-free algorithms to train this more complicated and bigger – compared to the simple feedforward models that had been experimented with in the literature so far – transformer architecture. It might be the case, that for large and complicated models it may be hard, or even almost impossible, to be trained from scratch using an evolutionary approach. Therefore, we wanted to experiment with first pretraining the agent using a supervised learning of sequence prediction on data generated by a smaller, potentially weaker model, which can be trained using any arbitrary reinforcement learning method. We are not searching for use cases where we gain an edge over existing approaches by using Decision Transformers trained via evolution; rather we aim to show that evolution works well even on complex models, and thus we need not hesitate to use the complex models with, e.g., evolutionary reinforcement learning algorithms [12, 13], the hybrid state-of-the-art reinforcement learning algorithms that combine gradients and evolution, performing better than the two individually.
The main contribution of this paper is showing that evolution strategies are capable of training the transformer architecture in the reinforcement learning setting – even though for large models they require a huge number of CPUs to run efficiently – which allows us to harness their exploration power even for transformers. We do so using OpenAI-ES [8], essentially the simplest distributional evolution strategy possible – if it works well, all the more advanced algorithms can only improve its performance.
In the following section, we present the background for our experiments. In Section III, we introduce our experiments and present their results, whereas in Section IV, we discuss these results. Finally, we conclude the paper in the last section.
@@ II-A Evolution Strategies
Evolutionary algorithms are a large and quite a successful family of black-box derivative-free optimization algorithms. They consist of various metaheuristics inspired by nature. One subclass of such nature-inspired optimization methods are evolution strategies. Introduced as a tool for dealing with high-dimensional continuous-valued domains [3], evolution strategies work with a population of individuals (real-valued vectors). In each generation (iteration / population in the iteration), they derive a new set of individuals by somehow mutating (perturbating) the original population; the new set is then evaluated (with respect to a given objective function), and a new generation is formed based on these new evaluated individuals taking into account their fitness (objective function value).
In order to use an evolution strategy as a reinforcement learning algorithm, we use agents represented by a neural network, or, more specifically, by a real-valued vector of the network weights, as the individuals for the algorithm. Then the only thing needed is to set the fitness of each individual as the mean return, the mean cumulative reward of the individual agent from several episodes. Various evolution strategy algorithms were proposed for this purpose [4].
@@
However, evolution strategies also offer many advantages. Many of the evolution strategy algorithms are easily parallelizable, and much of the research in the area has been focused on this aspect, yielding us algorithms with a linear improvement of performance when more computing power is used [8]. And because evolution strategies are derivative-free algorithms, we can optimize not just classical smooth neural networks, but our models can even contain some discrete subfunctions or be otherwise non-differentiable.
In this paper, we will be working with OpenAI-ES [8], which is a representative of Natural evolution strategies [14]. The population is represented by a distribution over the agent’s (neural network’s) parameters, the distribution being a Gaussian, whose mean value is our current solution to the given problem and from which the offsprings are sampled each generation. These are then evaluated, and their evaluation is used to update the parameters (in our case only the mean value) of the distribution so that we obtain a higher expected fitness for the future samples. This update is performed using an approximation of natural gradient (whence the name Natural evolution strategies). In our case, this is obtained by renormalizing (rescaling) the update with respect to uncertainty. (In general, we would have to compute an inverse of a so-called Fisher information matrix, which would then be used to multiply the gradient estimate by. However, as demonstrated, e.g., in a paper by Pierrot et al. [15], in our case, when we derive the parameter updates from the Gaussian distribution with the same variance for each parameter, dividing by the variance (the rescaling w.r.t. uncertainty) achieves approximately the same result.) The algorithm is also designed in such a way that it is highly parallelizable with interprocess communication kept at bare minimum. It shares many similarities with DeepGA [16], a descendant of genetic algorithms, another branch of the family of evolutionary algorithms. The main difference is that DeepGA has explicit population instead of the implicit population coded by the distribution of OpenAI-ES. For more details or for a discussion of the design choices, we refer our readers to the original paper [8].
@@ II-B Transformers
The transformers interlace classical feedforward layers with self-attention layers. (In practice, there are also residual connections and layer normalizations to stabilize the training.) And it is the self-attention to which the transformers owe their success. For each element of an input sequence, the self-attention constructs a “key”, a “query”, and a “value” using fully connected layers of neurons. Then, to compute an $i$-th element of an output sequence, it takes a combination of all the values, each weighted proportionally to the product of the query belonging to the given ($i$-th) position and a key corresponding to the value – hence combining the information from the whole input sequence to produce every single element of the output sequence, as shown by the following equation.
@@
Its main idea is that we want the agent’s policy to produce an action based on not just the last observation, but on the whole history (or the part which squeezes into the context window) of past observations and undertaken actions. And to have some way to affect the agent’s performance, we add a conditioning on a return-to-go, which is a return we want to obtain from a given step until the end of the episode.
@@
II-C Our OpenAI-ES Implementation
Mainly due to different computational resources available to us – we needed to use the MPI framework instead of EC2 (AWS) – as well as the need to interchange the models being trained and the environments for the evaluation more easily and with a greater degree of freedom, we decided to create our own implementation of OpenAI-ES. [^1] Our code, together with all the data gathered during our experiments, can be found on the following link: https://github.com/Mafi412/Evolution-Strategies-and-Decision-Transformers Yet, in the process of creating this implementation of ours we noticed some inconsistencies between the original paper introducing the algorithm [8] and their provided code. We decided to adhere more to the paper when these dissimilarities occured and here we state our reasoning for the two most significant differences.
First and foremost, in the paper [8], the use of a weight decay is mentioned as a form of keeping the effect of newly added updates still significant enough. After every update, we decay the weights of our model, which is the distribution mean. Therefore, it would not happen that there could be weights grown to such sizes that a newly added noise practically does not change the function of the model. This reasoning tells us why we want to use preciselly the weight decay. Nevertheless, in the original implementation, L2-regularization is used instead. This may be all right for some optimizers, as for, e.g., SGD or SGD with momentum it is virtually the same as the weight decay (just rescaled by the learning rate). However, as mentioned in paper Decoupled Weight Decay Regularization [18], it is not the same, for example, for ADAM optimizer, which is nonetheless exactly the one used in the original experiments. Our implementation, on the other hand, remains true to the original paper and uses proper weight decay. Nevertheless, we employ SGD with momentum in our experiments, as it demonstrated superior efficacy in our setting during our experiments.
Second, during the update of the distribution parameters, a gradient estimate should be normalized with respect to the uncertainty using a division by the standard deviation, so that it becomes a natural gradient estimate. Despite that, in the original OpenAI code, this is not done. There, this is probably hidden in the value of a learning rate (which, as a result, differs from ours). However, then the learning rate and the noise deviation are unnecessarily coupled hyperparameters.
In addition to the implementation of OpenAI-ES, we made a few slight modifications even in the Decision Transformer code. The most significant one concerns keeping data from the past to be used in future inputs. In the original code, the authors do not clip sequences already longer than the context window; those are hence allowed to grow unconstrained in the memory. The new data are always just appended to the list, keeping the whole history, and the whole list is passed to an inference function, which only then clips the sequence to the required length (without modifying the data stored) and passes it to the transformer. This proved to be no problem in some environments, but for some reason leads to a severe slowdown in others. Thus, we keep the stored sequence cropped to match the context length in the corresponding version of Decision Transformer.
@@ III Experiments
From what we saw in the previous section, the evolution strategies are viable and well-functioning algorithms for solving the reinforcement learning problem with promising parallelization and exploration abilities [19, 20], whereas the transformers embody a new possibility for how to approach this problem, as well as a potentially strong architecture of the neural network representing the agent’s policy. Therefore, it is just a logical next step to try and combine these two concepts, to test the capability of the transformers to be trained by evolution strategies in the setting of reinforcement learning.
As for the evolution strategy, we chose to use the OpenAI-ES since in its core it is the simplest evolution strategy possible. It uses a simple Gaussian distribution with its parameter covariance matrix being just a $\sigma\cdot I$, a rescaled identity matrix, hence no covariance is captured in this distribution, nor any difference in significance (or sensitivity) of different network weights. It does not even update the value of $\sigma$ hyperparameter during training. Therefore, if this evolution strategy works well, the others can only improve its performance.
@@
We provided our own implementation of the OpenAI-ES and began by a replication experiment of the original paper, which also serves as a correctness check for our code. Here, we tested the algorithm on a classical feedforward network in the MuJoCo [21] Humanoid environment using OpenAI Gym [22], whereupon we followed up by proceeding to test the performance of the evolution strategy with the Decision Transformer in the same environment. We then examined an idea of pretraining the transformer by a behavior cloning of some smaller and easily trained – yet possibly weaker – model before training the transformer by the evolution strategy. The last experiments we present are studying the particulars of training an even larger model in a different environment with a higher-dimensional image input, which is Hero, one of the Atari games [23].
Because training larger models takes more time, and so training the Decision Transformer is particularly time-intensive, we chose the Humanoid environment as a representative of MuJoCo environments, since it is the most complex and challenging among the standard ones. As for the Atari, because there training takes even longer, we selected the Hero game as an example of a medium-difficulty Atari game. We use Atari mainly to demonstrate training behavior on an even more complex environment with a more complex visual input, rather than to show that we can solve all its games.
@@
In all the experiments with Decision Transformers, we used the same hyperparameter values for the model as used in the original paper [10] for the respective environments. For comparison, the model sizes are 166 144 parameters for the feedforward model from the original OpenAI-ES paper [8], 825 098 parameters for the Decision Transformer for MuJoCo Humanoid environment and 2 486 272 parameters for Atari games environment.
For all the experiments, we conducted ten runs of the training. For every run of each of the experiments, 300 workers were used utilizing 301 CPU cores (with one being a master handling synchronization, evaluation, and saving the agent). Given the computational complexity of the task and the fact that our goal is to demonstrate that the evolution is able to train the transformer, we did not wait for the run to converge. Instead, at the beginning, we assigned each run a specific number of iterations it could use to train the model, and once these iterations were exhausted, we stopped the training. The desired returns passed to all the Decision Transformer models at the beginning of each episode were 7000 for the Humanoid Decision Transformer and 8000 for the Atari Decision Transformer. However, here we need to mention, that in the original Decision Transformer paper [10], for MuJoCo environments they rescale the rewards and returns-to-go dividing them by 1000, probably so that the values are somewhat normalized. For compatibility with the paper, we decided to keep this rescaling and hence the initial return-to-go value passed to the model when 7000 is passed as an argument to the algorithm is 7. All the figures reporting fitness progression during the evolution training in Humanoid environment work with these rescaled returns as well, because the rescaled values are those logged during the training. To obtain the fitness value, the return in the form the original Humanoid environment yields us, we only need to multiply the reported value by 1000.
@@
Figure 2: Median and quartiles of the best fitness values so far for all the main experiments done in the Humanoid environment for ten runs each. “FF” shows the training of a classical feedforward network using OpenAI-ES, as described in Section III-A. “DT” shows the training of a Decision Transformer using OpenAI-ES, as introduced in Section III-B. “Pre” shows the training of a pretrained Decision Transformer using OpenAI-ES, as discussed in Section III-C. (This experiment differs from the previous two by the fact that its number of iterations was smaller since the model was pretrained, which is the reason why its line ends in the middle of the plot.) Finally, the “TD3” horizontal line shows the final average performance of the best Decision Transformer trained by TD3 in a given time, as mentioned in the introduction of Section III.
To have a baseline to compare our experiments with the Decision Transformer to, we use the feedforward model trained using the OpenAI-ES during our first experiment described in Section III-A. However, to also have some initial comparison to the gradient-based methods, we wanted to try and train the Decision Transformer using gradients. The problem with this is that the Decision Transformer is originally intended as an offline reinforcement learning model. There exists a so-called Online Decision Transformer [24], but even it assumes starting with some pretraining. Although its architecture slightly differs from the classical Decision Transformer – it mainly differs in the action decoding layer, where a stochasticity is added – we tried to obtain the gradient baseline by training it using its standard training loop, just without any initial pretraining on precollected trajectories. However, the Online Decision Transformer proved to be unable to efficiently improve. Hence, we moved our attention to classical online reinforcement algorithms, even though those required a certain degree of creativity to make them work, since the Decision Transformer does not take only the last state as an input, but whole sequences of recent states, actions and the corresponing returns-to-go. Thus, we utilized a TD3 [25], a state of the art gradient reinforcement learning algorithm, implemented by Stable Baselines3 library [26]. We ran the training for 1 500 000 timesteps – with this number of timesteps on one CPU (environment simulation) and one GPU (gradient training) it ran for roughly a similar wall-clock time as our experiments with OpenAI-ES in Section III-B. Still, with more timesteps, the model would probably continue to improve. We conducted five runs of this gradient training in Humanoid environment and selected the final average performance of the best trained model as a baseline that is then shown in Figure 2. As we can see, the evolution is currently the only standard online reinforcement learning approach that can be applied to training the Decision Transformer without any adjustment, the gradient-based ones are either non-functional or need to be modified first.
III-A Feedforward - Humanoid